# Detecting Modeling Bias with Continuous Time Flow Models on Weak Lensing Maps

### v1

@@ Abstract
Simulation-based inference provides a powerful framework for extracting rich information from nonlinear scales in current and upcoming cosmological surveys, and ensuring its robustness requires stringent validation of forward models. In this work, we recast forward model validation as an out-of-distribution (OoD) detection problem within the framework of machine learning (ML)-based simulation-based inference (SBI). We employ probability density as the metric for OoD detection, and compare various density estimation techniques, demonstrating that field-level probability density estimation via continuous time flow models (CTFM) significantly outperforms feature-level approaches that combine scattering transform (ST) or convolutional neural networks (CNN) with normalizing flows (NFs), as well as NF-based field-level estimators, as quantified by the area under the receiver operating characteristic curve (AUROC). Our analysis shows that CTFM not only excels in detecting OoD samples but also provides a robust metric for model selection. Additionally, we verified CTFM maintains consistent efficacy across different cosmologies while mitigating the inductive biases inherent in NF architectures. Although our proof-of-concept study employs simplified forward modeling and noise settings, our framework establishes a promising pathway for identifying unknown systematics in the cosmology datasets.
@@ 1 Introduction
Beyond their impact on inference, these biases also signal gaps in our understanding of the underlying physics. Discrepancies between forward models and observations may arise from diverse sources—including cosmological evolution, complex astrophysical processes, and unaccounted observational effects, and thus merit thorough investigation. Detecting such biases can be framed as a consistency test: given that the full range of forward model outputs forms a statistical distribution, the task becomes one of determining whether an observation is a member of that distribution, which is often called the out-of-distribution (OoD) detection. When the likelihood $\mathcal{L}$ is Gaussian, $-2\log\mathcal{L}$ follows the $\chi^{2}$-distribution and is widely used to assess how well the model fits the data in current survey analysis [e.g. 39, 40, 41]. Recently, [38] generalizes the test to high-dimensional field-level inference where the likelihood is no longer Gaussian. They employed wavelet decomposition to segregate information by scale, leveraging the relative robustness of large-scale structures in modeling, and used normalizing flows to learn the corresponding distributions at each scale. Moreover, continuous time flow models (CTFM), such as diffusion models [42, 43, 44] and flow matching (FM) models [45], have emerged as state-of-the-art techniques for learning high-dimensional distributions. These methods have seen extensive application in the generation and inference of various cosmological fields [e.g. 46, 47, 48, 49, 50]. In this work, we explore the use of CTFM for bias detection at the field level and compare its performance against normalizing flows applied at both the field and feature levels.
@@ 2.2.1 Feature-level density estimation
In the feature-level density estimation, the high-dimensional field is first compressed into a low-dimensional feature vector, after which its probability density is estimated using a NF. Specifically, we employ two compressors: the Scattering Transform (ST) coefficients and a CNN trained with the VMIM loss [52], which provide optimal performance for ad-hoc and learned summary statistics, respectively. The real-valued non-volume preserving transformations (RealNVP) [53] is then used as the NF density estimator.
@@
$S_{1}(j,l)$ =$\bm{I}_{1}(j,l)\right>,$
$S_{2}(j_{1},l_{1},j_{2},l_{2})$ =$\bm{I}_{2}(j_{1},l_{1},j_{2},l_{2})\right>.$
[…]
@@ 2.2.2 Field-level density estimation
[…]
@@ 2.3.1 Dark matter only maps
The dark-matter-only (DMO) weak lensing (WL) convergence maps used in this study are obtained from [28] and are generated from a suite of 80 N-body simulations under flat $\Lambda$CDM cosmologies. In these simulations, the baryon density, Hubble parameter, and spectral index are fixed at $\Omega_{b}=0.046$, $h=0.72$, and $n_{s}=0.96$, respectively, while $\Omega_{m}$ and $\sigma_{8}$ vary around $\Omega_{m}=0.26$ and $\sigma_{8}=0.8$.
The N-body simulations employ $512^{3}$ particles in a box of size $240h^{-1}$ Mpc and are run using GADGET-2 [63]. WL convergence maps are computed via ray-tracing using a multiple-lens-plane algorithm [64] on snapshots spanning $10<z<$ 1. From each snapshot, an $80h^{-1}$ Mpc slice along the line-of-sight is extracted as a lens plane. By applying random rotations, flips, and shifts to the snapshots, 512 pseudo-independent maps are derived from each simulation. Further details on the map generation process can be found in [28].
@@ 3.1 Testing the field-level density estimator on Gaussian random fields
Gaussian random fields (GRFs) provide an ideal benchmark because their log–likelihood can be computed in closed form. Therefore, we first validates our density estimation formalism on GRFs. We generate GRFs whose power spectrum is a power law with additive white noise,
@@
and test three noise amplitudes, $\sigma_{n}\in\{1,2,3\}$. For each $\sigma_{n}$ we draw $120{,}000$ realisations of size $64^{2}$ pixels and train an OTFM model; its performance is assessed on an independent set of $1024$ samples, yielding the OTFM estimate $\log p_{\mathrm{OTFM}}$.
@@ 3.2 Detecting BCM maps as OoD
In this subsection, we present main results from our test problem, where we try to detect the unmodeled baryonic effects when DMO simulations are available. We test the performance of each methods mentioned above on our fiducial cosmology $(\sigma_{8},\Omega_{m})=(0.82,0.265)$. Unless stated otherwise, the InD samples consist of 304 maps of DMO maps at fiducial cosmology, while OoD samples are 1024 BCM maps with different baryon parameters at the same fiducial cosmology. This setting corresponding to an idealized $\delta$-function like posterior, and the reported performance here is the upper limit, as a broadened posterior leads to wider range of distribution of test statistics, therefore worse performance. These maps have a physical resolution of $128^{2}$ pixels. For each methods mentioned before, we compute the probability density of these maps as the $t(\bm{x})$, because the density is the most natural OoD detector. Examples with noise level $n_{g}=30$ are shown in Figure 3, and a well-behaved OoD detector is expected to generate clearly separable density distributions for InD and OoD sets.
@@
The Area Under the ROC Curve (AUROC) summarizes this plot into a single number ranging from 0 to 1. A larger AUROC indicates better overall classifier performance, with 0.5 representing a random guess and 1.0 indicating a perfect model. This metric is particularly useful for comparing different , as it remains invariant to class distribution and threshold selection.
@@ 3.4 OoD as a model selection
To validate the performance of generative model likelihoods as model selectors, we test the OTFM model in a noise miscalibration scenario, as OTFM is the best test statistics according to the results in Section 3. Specifically, we consider the mock observation $\bm{x}_{\rm mock}$, which consists of 304 DMO maps with shape noise $n_{g,{\rm ture}}=30{\rm arcmin^{-2}}$ at fiducial cosmology. Our three candidate models are DMO simulations with $n_{g}=\{25,30,35\}{\rm arcmin^{-2}}$, and their final noise standard deviations are $\sigma_{n}\approx\{0.0345,0.0315,0.0291\}$. We train an OTFM model for each candidate model $\mathcal{M}$ to serve as the density estimator $\log p(\bm{x}|\mathcal{M})$.
We first calculated the log density of $\bm{x}_{\rm mock}$ with different models and find their the log density distribution overlaps with each other as is shown in the left panel of Figure 5, making it is hard to identify the best model. However, the log density distribution of $\bm{x}_{\rm mock}$ for mis-specified models significantly differ from that of InD samples, shed a light to model comparison with the difference to the typical values of the InD samples. In this case, we can select the best model by the false positive rate of identify them as OoD, or equivalently the log density deviation from the log density of typical set samples. Therefore, the selection metric is then defined as
@@
where $p_{\rm InD}(x)$ is approximated by a dataset of model output. Intuitively, $t$ grows if (i) the model is overly broad—so its typical‑set likelihood is lower than that of the $\bm{x}_{\text{mock}}$—or (ii) the model assigns substantially lower likelihood to $\bm{x}_{\text{mock}}$ than to its own typical samples. The right panel of Figure 5 shows the distribution of $t(\bm{x}_{\rm mock}|\mathcal{M})$ for different $n_{g}$ models, demonstrating that the correct $n_{g}$ model attains a significantly smaller $t(\bm{x}|\mathcal{M})$. With this metric OTFM successfully selects the correct model for all 304 $\bm{x}_{\rm mock}$ maps.
@@ 3.5 Impact of resolution and survey area
Each of these OoD sets contains 304 samples. We then employed both the GLOW and OTFM models to estimate the probability density as test statistics $t(x)$; the corresponding PDFs are shown in Figure 8. It is evident that GLOW misestimates the OoD sets, whereas OTFM remains robust and successfully identifies all OoD samples. This limitation originates from the inductive bias inherent in the GLOW architecture, and the loss is less important than the architecture [70]. Although alternative test statistics, such as the 2-norm of the score function $\lVert\partial\log p(\bm{x})/\partial\bm{x}\rVert$ which serves as a metric for typicality, can mitigate this issue [71, 72], they still suffer from certain biases [73]. In conclusion, our findings confirm that CTFM generally outperforms NF, as it has less inductive biases from the model architecture, thus, the model can learn more information from the loss function without being limited by the model inductive bias.
@@ Acknowledgments
KD is supported by the National SKA Program of China (grant No. 2020SKA0110401) and NSFC (grant No. 11821303). BD acknowledges support from the Ambrose Monell Foundation, the Corning Glass Works Foundation Fellowship Fund, and the Institute for Advanced Study. This work is also supported by U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research under Contract No. DE-AC02-05CH11231 at Lawrence Berkeley National Laboratory to enable research for Data-intensive Machine Learning and Analysis. KD thanks Tao Jing, Ce Sui, Bhuvnesh Jain, Matias Zaldarriaga, Richard Grumitt and Xiaosheng Zhao for for their helpful insights and comments.
@@ A.1 CNN feature compressor
The 16 residual blocks are sequentially grouped into 4 groups with $\{3,4,6,3\}$ blocks, each group has $\{64,128,256,512\}$ channels, after each group, the map is 2×2$\times$2 × downsampled. Finally, the map is average-pooled to 512 feature maps and pass through a fully-connected layer to 128 output features.
@@ A.3 U-Net in CTFM
Apart from the input and output layers, our U-Net comprises 36 residual blocks in the encoder and an equal number in the decoder. Every three blocks, the feature map is downsampled by a factor of 2×2$\times$2 × in the encoder and upsampled by a factor of 2×2$\times$2 × in the decoder. When the map’s resolution falls below $16^{2}$, the middle block within each group of three at the same resolution is augmented with the attention mechanism. In this module, four distinct sets of $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are computed, constituting a so-called four-head attention, and the four $\mathbf{Y}$s are subsequently convolved to produce the final residual function learned by the block. This U-Net configuration is consistently applied in both the diffusion model and the OTFM across varying resolutions.

### v2

@@ Abstract
Simulation-based inference (SBI) provides a powerful framework for extracting rich information from nonlinear scales in current and upcoming cosmological surveys, and ensuring its robustness requires stringent validation of forward models. In this work, we recast forward model validation as an out-of-distribution (OoD) detection problem within the framework of machine learning (ML)-based SBI. We employ probability density as the metric for OoD detection, and compare various density estimation techniques, demonstrating that field-level probability density estimation via continuous time flow models (CTFM) significantly outperforms feature-level approaches that combine scattering transform (ST) or convolutional neural networks (CNN) with normalizing flows (NFs), as well as NF-based field-level estimators, as quantified by the area under the receiver operating characteristic curve (AUROC). Our analysis shows that CTFM not only excels in detecting OoD samples but also provides a robust metric for model selection. Additionally, we verified CTFM maintains consistent efficacy across different cosmologies while mitigating the inductive biases inherent in NF architectures. Although our proof-of-concept study employs simplified forward modeling and noise settings, our framework establishes a promising pathway for identifying unknown systematics in the cosmology datasets.
@@ 1 Introduction
Beyond their impact on inference, these biases also signal gaps in our understanding of the underlying physics. Discrepancies between forward models and observations may arise from diverse sources—including cosmological evolution, complex astrophysical processes, and unaccounted observational effects, and thus merit thorough investigation. Detecting such biases can be framed as a consistency test: given that the full range of forward model outputs forms a statistical distribution, the task becomes one of determining whether an observation is a member of that distribution, which is often called the out-of-distribution (OoD) detection. When the likelihood $\mathcal{L}$ is Gaussian, $2\log\mathcal{L}$ follows the $\chi^{2}$-distribution and is widely used to assess how well the model fits the data in current survey analysis [e.g. 39, 40, 41]. Recently, [38] generalizes the test to high-dimensional field-level inference where the likelihood is no longer Gaussian. They employed wavelet decomposition to segregate information by scale, leveraging the relative robustness of large-scale structures in modeling, and used normalizing flows to learn the corresponding distributions at each scale. Moreover, continuous time flow models (CTFM), such as diffusion models [42, 43, 44] and flow matching (FM) models [45], have emerged as state-of-the-art techniques for learning high-dimensional distributions. These methods have seen extensive application in sampling various cosmological fields across different tasks, including emulation [46], super-resolution [47] and reconstruction of cosmological fields[48, 49], but they are not yet widely used for computing the probability at the field level. [50] estimates the lower bound of field-level probability with CTFM, while in this work we compute the field-level probability directly through integral of CTFM trajectory. We explore the use of CTFM probability density for bias detection at the field level and compare its performance against normalizing flows applied at both the field and feature levels.
@@ 2.2.1 Feature-level density estimation
In the feature-level density estimation, the high-dimensional field is first compressed into a low-dimensional feature vector, after which its probability density is estimated using a NF. In this work, we employ two compressors: the Scattering Transform (ST) coefficients and a CNN trained with the VMIM loss [52]. Specifically, the CNN-learned statistic has been shown to provide optimal performance in constraining cosmological parameters on Gaussian random fields and log-normal fields [53, 54], while on non-linear weak lensing maps, it also achieves similar performance to field-level analysis [55]. After the compression, the real-valued non-volume preserving transformations (RealNVP) [56] is then used as the NF density estimator.
@@
$S_{1}(j,l)$ $=\left<\bm{I}_{1}(j,l)\right$>,
$S_{2}(j_{1},l_{1},j_{2},l_{2})$ $=\left<\bm{I}_{2}(j_{1},l_{1},j_{2},l_{2})\right$>.
Here, $\bm{x}$ is the input field, $\ast$ denotes the convolution operation, $\Psi$ represents the Morlet wavelet kernel (see e.g. Appendix B of [25] for details) and $\Phi$ is the Gaussian kernel to filter all small-scale fluctuations. The index $j$ specifies the scale of the convolutional kernel with smaller $j$ corresponding to more localized kernels while $l$ defines its orientation. We select $j=0\text{–}3$ and $l=0\text{–}3$ to cover a broad range of scales and orientations, yielding 16 coefficients for $S_{1}(j,l)$ and 96 coefficients for $S_{2}(j_{1},l_{1},j_{2},l_{2})$, for a total feature vector length of 112. The ST coefficients are computed using Kymatio [^1] https://github.com/kymatio/kymatio [57].
@@ 2.2.2 Field-level density estimation
with $\alpha_{t}:=\exp\left(-\frac{1}{2}\int_{0}^{t}\beta_{s}\,ds\right)$. Since the probability flow ODE and the SDE share the same marginal distribution $p_{t}(\bm{x}_{t})$, they also share the same score function. By substituting the neural network $\bm{s}_{\bm{w}}(\bm{x},t)$ in place of the true score in Equation 2.11, we obtain the diffusion model ODE for estimating $p(\bm{x})$. Our implementation of the diffusion model, including the neural network, is based on the diffusers [^2] https://huggingface.co/docs/diffusers/index package. We choose the linear schedule $\beta_{t}=20t$ in this work.
@@ 2.3.1 Dark matter only maps
[…]
The N-body simulations employ $512^{3}$ particles in a box of size $240h^{-1}$ Mpc and are run using GADGET-2 [66]. WL convergence maps are computed via ray-tracing using a multiple-lens-plane algorithm [67] on snapshots spanning $0<z<1$. From each snapshot, an $80h^{-1}$ Mpc slice along the line-of-sight is extracted as a lens plane. By applying random rotations, flips, and shifts to the snapshots, 512 pseudo-independent maps are derived from each simulation. Further details on the map generation process can be found in [28].
@@ 3.1 Testing the field-level density estimator on Gaussian random fields
Gaussian random fields (GRFs) provide an ideal benchmark for probability density estimation methods because their log–likelihood can be computed in closed form. Therefore, we first validates our density estimation formalism on GRFs. We generate GRFs whose power spectrum is a power law with additive white noise,
@@
and test three noise amplitudes, $\sigma_{n}\in\{1,2,3\}$. For each $\sigma_{n}$, we draw 120,000120{,}000 realisations of size $64^{2}$ pixels and train an OTFM model; its performance is assessed on an independent set of $1024$ samples, yielding the OTFM estimated $\log p_{\mathrm{OTFM}}$ by applying OTFM to Equation 2.8.
@@
While we do not explicitly test OoD detection here, the accurate density estimation suggests OTFM’s potential for identifying OoD samples in such GRF datasets. We also expect CNN and ST statistics to achieve good performance on this simple dataset. The former has been shown to extract the full information contents on GRFs and lognormal fields [53, 54], while the first-order ST coefficients are strongly correlated with the power spectrum [25].
@@ 3.2 Detecting BCM maps as OoD
In this subsection, we present main results from our test problem, where we try to detect the unmodeled baryonic effects when DMO simulations are available. We test the performance of each methods mentioned above on our fiducial cosmology $\theta_{\rm fid}=(\sigma_{8,\rm fid},\Omega_{m,\rm fid})=(0.82,0.265)$. For feature-level detection methods like CNN and ST, the InD samples consist of 304 maps of DMO maps at fiducial cosmology, while OoD samples are 1024 BCM maps with different baryon parameters at the same fiducial cosmology. This setting corresponds to an idealized case where we can perfectly infer the cosmological parameters, i.e. $p(\theta|\bm{x}_{\rm obs})=\delta(\theta-\theta_{\rm fid})$ in Equation 2.1, to avoid sampling the full posterior which is computationlly expensive. Using the true $p(\theta|\bm{x}_{\rm obs})$ will introduce $\mathcal{O}(1)$ scatter to the test statistics and slightly degraded OoD performance. Thus our reported CNN and ST performance here represent the upper limit. These maps have a physical resolution of $128^{2}$ pixels. For each methods mentioned before, we compute the probability density of these maps as the $t(\bm{x})$ using Equation 2.8, because the density is the most natural OoD detector. Examples with noise level $n_{g}=30$ are shown in Figure 3, and a well-behaved OoD detector is expected to generate clearly separable density distributions for InD and OoD sets.
@@
The Area Under the ROC Curve (AUROC) summarizes this plot into a single number ranging from 0 to 1. A larger AUROC indicates better overall classifier performance, with 0.5 representing a random guess and 1.0 indicating a perfect model. This metric is particularly useful for comparing different classifiers, as it remains invariant to class distribution and threshold selection.
@@
When the type of OoD is known (i.e., OoD data is given at training stage), one can use supervised training techniques to train CNNs and obtain better performance [e.g. 70, 71]. However, in many applications one does not know the type of OoD, and instead tries to search for unknown unknowns. This is the main focus of our paper, and supervised training is not applicable here since no OoD training data is given. In this scenario, we expect training the CNN with an AutoEncoder (AE) loss could improve OoD detection compared to the VMIM loss we tried, as the AE’s reconstruction task necessitates learning a comprehensive representation of the input field, rather than just cosmology-specific features, and we plan to explore this in our future works.
@@ 3.4 OoD as a model selection
To validate the performance of generative model likelihoods as model selectors, we test the OTFM model in a noise miscalibration scenario, as OTFM is the best test statistics according to the results in Section 3. Specifically, we consider the mock observation $\bm{x}_{\rm mock}$, which consists of 304 DMO maps with shape noise $n_{g,{\rm{true}}}=30{\rm arcmin^{-2}}$ at fiducial cosmology. Our three candidate models are DMO simulations with $n_{g}=\{25,30,35\}{\rm arcmin^{-2}}$, and their final noise standard deviations are $\sigma_{n}\approx\{0.0345,0.0315,0.0291\}$. We train an OTFM model for each candidate model $\mathcal{M}$ to serve as the density estimator $\log p(\bm{x}|\mathcal{M})$.
We first calculated the log density of $\bm{x}_{\rm mock}$ with different models and find their the log density distribution overlaps with each other as is shown in the left panel of Figure 5, making it hard to identify the best model. However, the log density distribution of $\bm{x}_{\rm mock}$ for mis-specified models significantly differ from that of InD samples, shed a light to model comparison with the difference to the typical values of the InD samples. In this case, we can select the best model by the false positive rate of identify them as OoD, or equivalently the log density deviation from the log density of typical set samples. Therefore, the selection metric is then defined as
@@
where $p_{\rm InD}(x)$ is approximated by a dataset of model output, $\mathbb{E}_{\bm{x}\sim p_{\rm InD}(\bm{x})}\bigl{(}\log p(\bm{x}|\mathcal{M})\bigr{)}$ is the expectation of log density over the InD distribution, and is estimated empirically by taking average log density of the InD samples. Intuitively, $t$ grows if (i) the model is overly broad—so its typical‑set likelihood is lower than that of the $\bm{x}_{\text{mock}}$—or (ii) the model assigns substantially lower likelihood to $\bm{x}_{\text{mock}}$ than to its own typical samples. The right panel of Figure 5 shows the distribution of $t(\bm{x}_{\rm mock}|\mathcal{M})$ for different $n_{g}$ models, demonstrating that the correct $n_{g}$ model attains a significantly smaller $t(\bm{x}|\mathcal{M})$. With this metric OTFM successfully selects the correct model for all 304 $\bm{x}_{\rm mock}$ maps.
@@ 3.5 Impact of resolution and survey area
Each of these OoD sets contains 304 samples. We then employed both the GLOW and OTFM models to estimate the probability density as test statistics $t(x)$; the corresponding PDFs are shown in Figure 8. An ideal likelihood estimator is supposed to assign lower likelihood to these artificial OoD maps, however it is evident that GLOW misestimates the OoD sets by assigning higher likelihood for OoD samples, as is shown in the left panel of Figure 8. Meanwhile, OTFM remains robust and successfully identifies all OoD samples by assigning low likelihoods for these samples. This limitation originates from the inductive bias inherent in the GLOW architecture, and the loss is less important than the architecture [74]. Although alternative test statistics, such as the 2-norm of the score function $\lVert\partial\log p(\bm{x})/\partial\bm{x}\rVert$ which serves as a metric for typicality, can mitigate this issue [75, 76], they still suffer from certain biases [77]. In conclusion, our findings confirm that CTFM generally outperforms NF, as it has less inductive biases from the model architecture, thus, the model can learn more information from the loss function without being limited by the model inductive bias.
@@ Acknowledgments
KD is supported by the National SKA Program of China (grant No. 2020SKA0110401) and NSFC (grant No. 11821303). BD acknowledges support from the Ambrose Monell Foundation, the Corning Glass Works Foundation Fellowship Fund, and the Institute for Advanced Study. This work is also supported by U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research under Contract No. DE-AC02-05CH11231 at Lawrence Berkeley National Laboratory to enable research for Data-intensive Machine Learning and Analysis. KD thanks Tao Jing, Ce Sui, Bhuvnesh Jain, Matias Zaldarriaga, Richard Grumitt and Xiaosheng Zhao for their helpful insights and comments.
@@ A.1 CNN feature compressor
The 16 residual blocks are sequentially grouped into 4 groups with $\{3,4,6,3\}$ blocks, each group has $\{64,128,256,512\}$ channels, after each group, the map is $2\times$ downsampled. Finally, the map is average-pooled to 512 feature maps and pass through a fully-connected layer to 128 output features.
@@ A.3 U-Net in CTFM
Apart from the input and output layers, our U-Net comprises 36 residual blocks in the encoder and an equal number in the decoder. Every three blocks, the feature map is downsampled by a factor of $2\times$ in the encoder and upsampled by a factor of $2\times$ in the decoder. When the map’s resolution falls below $16^{2}$, the middle block within each group of three at the same resolution is augmented with the attention mechanism. In this module, four distinct sets of $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are computed, constituting a so-called four-head attention, and the four $\mathbf{Y}$s are subsequently convolved to produce the final residual function learned by the block. This U-Net configuration is consistently applied in both the diffusion model and the OTFM across varying resolutions.